OpenStudy (anonymous):

A bag contains 6 red marbles, 4 blue marbles, 7 green marbles, and 3 yellow marbles. A marble is drawn from the bag and is not replaced. Then a second marble is drawn. What is the probability that both marbles drawn were green? Option A: 49/380 Option B: 49/400 Option C: 21/190 Option D: 21/200 Please explain! I have no idea what it is and how to do it :(

OpenStudy (anonymous):

Is there some kind of equation i need to use to find the answer?? OMG I am soooo lost! D;

OpenStudy (cwtan):

\[\frac{7}{6+4+7+3}\times \frac {7-1}{6+4+7+3-1}\]

OpenStudy (anonymous):

So is the answer option C 21/190 then??

OpenStudy (cwtan):

If it is the answer of the equation i typed above then it is the answer~ Need my explaination?

OpenStudy (anonymous):

yes please! :) So the answer is 21/190 then??

OpenStudy (anonymous):

@cwtan is the answer 21/190 then??? I am getting kinda confused now ;O

OpenStudy (anonymous):

cuz i plugged your equation into a calculator and I got 0.11052, which is the same as 21/190 which is 0.11052... so C is the answer yes??!

OpenStudy (cwtan):

Yes Probability of a event to occur= Number of the event will be occur/Total Number of event So it is \[\frac {7}{20}\] It is being subtract one for numerator and denominator each because the number of green marbles is -1 and total has been -1.(not replaced)

OpenStudy (anonymous):

cool! thanks :)